# Coloring and density theorems for configurations of a given volume

### v2

@@ 4. Parallelograms: proof of Theorem 6
for integers $10\leqslant j,k\leqslant N-$ 1. For every monochromatic parallelogram $\mathcal{P}$ coordinatized as in Section 2 (see Figure 4) we conclude
@@
and we again conclude that this area cannot be equal to $1$. This is not surprising because we have in fact been working with the quantity $\mathscr{J}$ from (3.3) in Section 3 and we have just shown that it is invariant, but only for these special parallelograms. Rotate the coloring $\mathscr{C}^{\prime\prime}_{0},\ldots,\mathscr{C}^{\prime\prime}_{4}$ in a way that it prevents monochromatic area $1$ parallelograms from having one side parallel to one of the given lines $\ell_{i}$; repeat this for $i=1,2,\ldots,m$.
@@ 5. Hypercube graphs: proof of Theorem 8
for $10\leqslant j,k\leqslant 2^{n+1}-$ 1. If all $2^{n}$ points (1.3) belong to the same class $\mathscr{C}_{j,k}$, then (5.1) implies
@@ 6. Parallelotopes: proof of Theorem 7
By the induction on a positive integer $n$ we will prove that every $n$-parallelotope with vertices in the set
@@
The induction basis $n=1$ is trivial since the observed $1$-parallelotopes are just segments in the interval $(0,\theta]$, so their length is clearly less than $\theta$. Now take $n\geqslant 2$ and suppose that the claim holds for the sets $A_{n-1,\vartheta}$ for every $\vartheta>0$. Let $\theta>0$ and let $\mathcal{P}$ be an $n$-parallelotope with vertices in $A_{n,\theta}$. Suppose that ${p}$ is the vertex of $\mathcal{P}$ with the smallest first coordinate (i.e., coordinate $x_{1}$). Also suppose that $\mathcal{P}$ is spanned from ${p}$ by vectors ${v}_{1},\ldots,{v}_{n}$ and introduce their coordinates just as in formula (3.5). We have
@@
by the choice of the vertex ${p}$. Finally, for each $1\leqslant i\leqslant n$, let $\pi_{i}$ denote the hyperplane orthogonal to the first coordinate axis and passing through the point ${p}+{v}_{i}$ (i.e., with Cartesian equation $x_{1}=p_{1}+v_{i,1}$), let $\mathcal{P}_{i}$ denote the $(n-1)$-parallelotope spanned from the point ${p}+{v}_{i}$ by the vectors ${v}_{1},\ldots,{v}_{i-1},{v}_{i+1},\ldots,{v}_{n}$, and let $\mathcal{P}^{\prime}_{i}$ be the orthogonal projection of $\mathcal{P}_{i}$ onto $\pi_{i}$; see Figure 8. From (6.1) we see that all vertices of $\mathcal{P}_{i}$ have $x_{1}$-coordinates at least $p_{1}+v_{i,1}$, so the vertices of $\mathcal{P}^{\prime}_{i}$ lie in the set
@@ 7. Notation and preliminaries from harmonic analysis
From this section on we introduce the notation
@@
which means that inequality $\mathcal{A}\leqslant C_{P}\mathcal{B}$ holds with come constant $C_{P}\in(0,\infty)$ depending only on a set of parameters $P$. It is always understood that inequality constants depend on the (configuration or ambient) dimensions, so these are not explicitly listed in the set $P$. Next,
@@
Dimension $d$ is not visible from the notation $\mathbbm{g},\mathbbm{h}^{(l)},\mathbbm{k}$, but it will always be understood from a context.
@@
see [51, §VIII.3]. In particular, this holds when $\mu=\sigma$ is the spherical measure supported on the standard unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, also called the circle measure when $d=2$. Note that an easy consequence of (7.8) and (7.1) is
@@
when the exponents $p,q,r\in[1,\infty]$ satisfy $1+1/r=1/p+1/q$; see [52, §V.1]
@@ 8. Right simplices: proofs of Theorem 1 and Corollary 2
We only need to prove part (a) of Theorem 1. It is sufficient to work in dimension $n=m+1$, since, in higher dimensions $n$, Fubini’s theorem guaranties that at least one the sections parallel to a fixed $(m+1)$-dimensional coordinate plane has density at least that of $A$. The parameter $m$ is fixed from now on. Take an $R\in(0,\infty)$ and a measurable set $A\subseteq[0,R]^{m+1}$ with density
@@
where $\sigma$ is the circle measure in $\mathbb{R}^{2}$ normalized to have total mass $1$ on $\mathbb{S}^{1}$ and we write
@@
belong to $A$ by performing normalized integration over all simplex vertices $(x,y)\in[0,R]^{m+1}$, over $m-1$ simplex edges parallel to the coordinate axes with lengths $\theta\lambda\leqslant|u_{k}|\leqslant\lambda$, and over the last perpendicular edge parallel to the remaining coordinate $2$-plane with its length precisely equal to $|v|=m!|u_{1}\cdots u_{m-1}|^{-1}$, so that the $m$-dimensional simplex volume equals $1$. Merely knowing that $\mathcal{N}_{\lambda}^{0}(A;R)>0$ guarantees existence of the desired point configuration inside $A$. The freedom to choose the parameter $\lambda$ will allow certain pigeonholing. The idea is to show that $\mathcal{N}_{\lambda}^{0}$ is positive for at least one out of sufficiently many choices for $\lambda$, i.e., from a sufficiently large interval on the logarithmic scale, which is available as soon as $R$ is large enough (in terms of $\delta$).
@@
×$\times(\sigma\ast\mathbbm{g}_{\varepsilon})_{m!|u_{1}\cdots u_{m-1}|^{-1}}(v)\,\lambda^{-m+1}\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}v\,\textup{d}u\,R^{-m-1}\,\textup{d}y\,\textup{d}x.$
@@
×$\times R^{-m-1}\lambda^{-m+1}\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}u$
@@
where the names of the structured, error, and uniform parts are borrowed from regularity lemmas in extremal combinatorics, which they mimic. The structured part is a dominant and positive one, the uniform part is always small, while the error part can be made small only after successful pigeoholing in sufficiently many choices of $\lambda$; see a general discussion of the approach in [40, §1.3].
@@
$D:=\big{\{}(u,v,x,y):$ $|u_{k}|\leqslant\lambda\text{ for every }1\leqslant k\leqslant m-1,$
@@
Let $\mathcal{I}$ denote a partition of $[0,\lceil R/\lambda\rceil\lambda]\supseteq[0,R]$ into intervals of length $\lambda$ and let $\mathcal{Q}$ be a partition of $[0,\lceil R\lambda^{m-1}\rceil\lambda^{-m+1}]^{2}\supseteq[0,R]^{2}$ into squares of side-length $\lambda^{-m+1}$. Clearly, by the construction and the assumption $R\geqslant\max\{\lambda,\lambda^{-m+1}\}$,
@@
$\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\Big{(}\prod_{k=1}^{m-1}\int_{I_{k}}\mathbbm{1}_{A}(x_{1},\ldots,x_{k-1},x^{\prime},x_{k+1},\ldots,x_{m-1},y)\,\textup{d}x^{\prime}\Big{)}$
×$\times\Big{(}\int_{Q}\mathbbm{1}_{A}(x,y^{\prime})\,\textup{d}y\Big{)}\mathbbm{1}_{A}(x,y)\,\textup{d}(x,y)$ (8.7)
@@
By Hölder’s inequality, the integral (8.7) multiplied by
$\prod_{k=1}^{m-1}\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\frac{\mathbbm{1}_{A}(x,y)}{\int_{I_{k}}\mathbbm{1}_{A}(\ldots,x_{k-1},x^{\prime},x_{k+1},\ldots,y)\,\textup{d}x^{\prime}}\,\textup{d}(x,y)=\big{(}\lambda^{m-2}(\lambda^{-m+1})^{2}\big{)}^{m-1}$
(where the fraction is interpreted as 00 if the denominator equals 00) and
$\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\frac{\mathbbm{1}_{A}(x,y)}{\int_{Q}\mathbbm{1}_{A}(x,y^{\prime})\,\textup{d}y^{\prime}}\,\textup{d}(x,y)=\lambda^{m-1}$
is at least
$\Big{(}\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\mathbbm{1}_{A}(x,y)\,\textup{d}(x,y)\Big{)}^{m+1}=|A\cap(I_{1}\times\cdots\times I_{m-1}\times Q)|^{m+1}.$
@@
needs to be bounded from above. On the part of the set $E$ where $|u_{k}|<\theta\lambda$ for a fixed given index $11\leqslant k\leqslant m-$ 1 we simply estimate $\Phi$ by $1$ and integrate out in all variables, bounding that part of the integral by
@@
and only remember that $f(\cdot;u)$ is bounded by $1$ and supported on $[0,R]^{m+1}$. Also define
@@
×$\times(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}t}{2\pi t}$
@@
$\lesssim\int_{\tau}^{\varepsilon}t^{-1/2}\,\textup{d}t\lesssim\varepsilon^{1/2}$
@@
$R^{-m-1}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{2\pi t}$
@@
The corresponding integral of $1/s$ equals $1$, so the above equality is still valid. Whenever $\lambda\in(0,\infty)$, $t\in[\varepsilon,1]$, $s$ is as in (8.12), and $\theta\lambda\leqslant|u_{k}|\leqslant\lambda$ for $11\leqslant k\leqslant m-$ 1, then we define
@@
×$\qquad\qquad\qquad\qquad\times\Big{(}\int_{\mathbb{R}^{2}}\mathbbm{1}_{A}(x,y^{\prime})\mathbbm{h}^{(l)}_{s}(y^{\prime}-z^{\prime})\,\textup{d}y^{\prime}\Big{)}(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\,\textup{d}z\,\textup{d}z^{\prime}.$
@@
×$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}.$
@@
×$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}$
@@
×$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}$
@@
×$\times\int_{\varepsilon}^{1}\int_{\mathbb{R}}\int_{te^{-(m-1)\alpha-1}}^{te^{-(m-1)\alpha}}\int_{\mathbb{R}^{m-1}}\int_{\mathbb{R}^{2}}\big{|}\widehat{\mathbbm{1}_{A}}(x,\xi)\big{|}^{2}e^{-2\pi s^{2}|\xi|^{2}}|\xi|^{2}\,\textup{d}\xi\,\textup{d}x\,s\,\textup{d}s\,\textup{d}\alpha\,\frac{\textup{d}t}{t}.$
@@
where $C_{m}$ is the constant from Theorem 1. If we can find an axes-aligned cube $Q\subset\mathbb{R}^{n}$ of side-length $R_{0}$ on which $A$ has density greater than $(C_{m}/\log R_{0})^{1/(9m^{2})}$, then part (a) of that same theorem will apply to this cube $Q$, only translated from the position $[0,R_{0}]^{n}$. It will give us a right $m$-simplex of volume $1$ with vertices in $A\cap Q$. Moreover, the ratio of lengths of any two of its perpendicular edges will be at most
@@
Finally, for an arbitrary $V>0$ the set $V^{-1/m}A$ clearly has the same upper Banach density $\bar{\delta}_{n}(A)$, so the previous case applies to it and gives a right simplex of volume $1$ with vertices in $V^{-1/m}A$ and with controlled ratios of lengths of its perpendicular edges. It remains to scale the whole picture back by factor $V^{1/m}$ and thus obtain the desired simplex of volume $V$. ∎

### v3

@@ 4. Parallelograms: proof of Theorem 6
for integers $0\leqslant j,k\leqslant N-1$. For every monochromatic parallelogram $\mathcal{P}$ coordinatized as in Section 2 (see Figure 4) we conclude
@@
and we again conclude that this area cannot be equal to $1$. This is not surprising because we have in fact been working with the quantity $\mathscr{J}$ from (3.3) in Section 3 and we have just shown that it is invariant, but only for these special parallelograms. Rotate the coloring $\mathscr{C}^{\prime\prime}_{0},\ldots,\mathscr{C}^{\prime\prime}_{4}$ in such a way that it prevents monochromatic area $1$ parallelograms from having one side parallel to one of the given lines $\ell_{i}$; repeat this for $i=1,2,\ldots,m$.
@@ 5. Hypercube graphs: proof of Theorem 8
for $0\leqslant j,k\leqslant 2^{n+1}-1$. If all $2^{n}$ points (1.3) belong to the same class $\mathscr{C}_{j,k}$, then (5.1) implies
@@ 6. Parallelotopes: proof of Theorem 7
By induction on a positive integer $n$ we will prove that every $n$-parallelotope with vertices in the set
@@
The induction basis $n=1$ is trivial since the observed $1$-parallelotopes are just segments in the interval $(0,\theta]$, so their length is clearly less than $\theta$. Now take $n\geqslant 2$ and suppose that the claim holds for the sets $A_{n-1,\vartheta}$ for every $\vartheta>0$. Let $\theta>0$ and let $\mathcal{P}$ be an $n$-parallelotope with vertices in $A_{n,\theta}$. Suppose that p{p} is the vertex of $\mathcal{P}$ with the smallest first coordinate (i.e., coordinate $x_{1}$). Also suppose that $\mathcal{P}$ is spanned from p{p} by vectors ${v}_{1},\ldots,{v}_{n}$ and introduce their coordinates just as in formula (3.5). We have
@@
by the choice of the vertex p{p}. Finally, for each $1\leqslant i\leqslant n$, let $\pi_{i}$ denote the hyperplane orthogonal to the first coordinate axis and passing through the point ${p}+{v}_{i}$ (i.e., with Cartesian equation $x_{1}=p_{1}+v_{i,1}$), let $\mathcal{P}_{i}$ denote the (n1)(n-1)-parallelotope spanned from the point ${p}+{v}_{i}$ by the vectors ${v}_{1},\ldots,{v}_{i-1},{v}_{i+1},\ldots,{v}_{n}$, and let $\mathcal{P}^{\prime}_{i}$ be the orthogonal projection of $\mathcal{P}_{i}$ onto $\pi_{i}$; see Figure 8. From (6.1) we see that all vertices of $\mathcal{P}_{i}$ have $x_{1}$-coordinates at least $p_{1}+v_{i,1}$, so the vertices of $\mathcal{P}^{\prime}_{i}$ lie in the set
@@ 7. Notation and preliminaries from harmonic analysis
From this section on, we introduce the notation
@@
which means that inequality $\mathcal{A}\leqslant C_{P}\mathcal{B}$ holds with some constant $C_{P}\in(0,\infty)$ depending only on a set of parameters $P$. It is always understood that inequality constants depend on the (configuration or ambient) dimensions, so these are not explicitly listed in the set $P$. Next,
@@
Dimension $d$ is not visible from the notation $\mathbbm{g},\mathbbm{h}^{(l)},\mathbbm{k}$, but it will always be understood from context.
@@
see [56, §VIII.3]. In particular, this holds when $\mu=\sigma$ is the spherical measure supported on the standard unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, also called the circle measure when $d=2$. Note that an easy consequence of (7.8) and (7.1) is
@@
when the exponents $p,q,r\in[1,\infty]$ satisfy $1+1/r=1/p+1/q$; see [57, §V.1]
@@ 8. Right simplices: proofs of Theorem 1 and Corollary 2
We only need to prove part (a) of Theorem 1. It is sufficient to work in dimension $n=m+1$, since, in higher dimensions $n$, Fubini’s theorem guarantees that at least one of the sections parallel to a fixed $(m+1)$-dimensional coordinate plane has density at least that of $A$. The parameter $m$ is fixed from now on. Take an $R\in(0,\infty)$ and a measurable set $A\subseteq[0,R]^{m+1}$ with density
@@
where $\sigma$ is the circle measure in $\mathbb{R}^{2}$ normalized to have total mass 11 on $\mathbb{S}^{1}$ and we write
@@
belong to $A$ by performing normalized integration over all simplex vertices $(x,y)\in[0,R]^{m+1}$, over m1m-1 simplex edges parallel to the coordinate axes with lengths $\theta\lambda\leqslant|u_{k}|\leqslant\lambda$, and over the last perpendicular edge parallel to the remaining coordinate $2$-plane with its length precisely equal to $|v|=m!|u_{1}\cdots u_{m-1}|^{-1}$, so that the $m$-dimensional simplex volume equals $1$. Merely knowing that $\mathcal{N}_{\lambda}^{0}(A;R)>0$ guarantees existence of the desired point configuration inside $A$. The freedom to choose the parameter $\lambda$ will allow certain pigeonholing. The idea is to show that $\mathcal{N}_{\lambda}^{0}$ is positive for at least one out of sufficiently many choices for $\lambda$, i.e., from a sufficiently large interval on the logarithmic scale, which is available as soon as $R$ is large enough (in terms of $\delta$).
@@
$\times(\sigma\ast\mathbbm{g}_{\varepsilon})_{m!|u_{1}\cdots u_{m-1}|^{-1}}(v)\,\lambda^{-m+1}\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}v\,\textup{d}u\,R^{-m-1}\,\textup{d}y\,\textup{d}x.$
@@
$\times R^{-m-1}\lambda^{-m+1}\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}u$ .
@@
where the names of the structured, error, and uniform parts are borrowed from regularity lemmas in extremal combinatorics, which they mimic. The structured part is the dominant and positive one, the uniform part is always small, while the error part can be made small only after successful pigeonholing in sufficiently many choices of $\lambda$; see a general discussion of the approach in [43, §1.3].
@@
$D:=\big\{(u,v,x,y):\$$|u_{k}|\leqslant\lambda\text{ for every }1\leqslant k\leqslant m-1,$
@@
Let $\mathcal{I}$ denote a partition of $0,\lceil R/\lambda\rceil\lambda]\supseteq[0,R]$ into intervals of length $\lambda$ and let $\mathcal{Q}$ be a partition of $0,\lceil R\lambda^{m-1}\rceil\lambda^{-m+1}]^{2}\supseteq[0,R]^{2}$ into squares of side-length $\lambda^{-m+1}$. Clearly, by the construction and the assumption $R\geqslant\max\{\lambda,\lambda^{-m+1}\}$,
@@
$\mathcal{A}:=\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\Big(\prod_{k=1}^{m-1}\int_{I_{k}}\mathbbm{1}_{A}(x_{1},\ldots,x_{k-1},x^{\prime},x_{k+1},\ldots,x_{m-1},y)\,\textup{d}x^{\prime}\Big)$
$\times\Big(\int_{Q}\mathbbm{1}_{A}(x,y^{\prime})\,\textup{d}y\Big)\mathbbm{1}_{A}(x,y)\,\textup{d}(x,y)$ (8.7)
@@
Observe the equalities
$\mathcal{B}$ $:=\prod_{k=1}^{m-1}\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\frac{\mathbbm{1}_{A}(x,y)}{\int_{I_{k}}\mathbbm{1}_{A}(\ldots,x_{k-1},x^{\prime},x_{k+1},\ldots,y)\,\textup{d}x^{\prime}}\,\textup{d}(x,y)$
$\,=\big(\lambda^{m-2}(\lambda^{-m+1})^{2}\big)^{m-1}=\lambda^{-(m-1)m}$
and
$\mathcal{C}:=\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\frac{\mathbbm{1}_{A}(x,y)}{\int_{Q}\mathbbm{1}_{A}(x,y^{\prime})\,\textup{d}y^{\prime}}\,\textup{d}(x,y)=\lambda^{m-1},$
where the fraction $0/0$ is interpreted as 0. We now apply Hölder’s inequality for integration over $I_{1}\times\cdots\times I_{m-1}\times Q$, with $m+1$ factors and all exponents equal $m+1$. It gives
$\mathcal{A}\,\mathcal{B}\,\mathcal{C}\geqslant\Big(\int_{I_{1}\times\cdots\times I_{m-1}\times Q}\mathbbm{1}_{A}(x,y)^{m+1}\,\textup{d}(x,y)\Big)^{m+1},$
so the integral $\mathcal{A}$ from (8.7) is at least
$\lambda^{(m-1)^{2}}|A\cap(I_{1}\times\cdots\times I_{m-1}\times Q)|^{m+1}.$
@@
needs to be bounded from above. On the part of the set $E$ where $|u_{k}|<\theta\lambda$ for a fixed given index $1\leqslant k\leqslant m-1$ we simply estimate $\Phi$ by $1$ and integrate out in all variables, bounding that part of the integral by
@@
and only remember that $f(\cdot;u)$ is bounded by $1$ and supported on $0,R]^{m+1}$. Also define
@@
$\times(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}t}{2\pi t}$ .
@@
$\lesssim\int_{\tau}^{\varepsilon}t^{-1/2}\,\textup{d}t\lesssim\varepsilon^{1/2}$ .
@@
$R^{-m-1}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{2\pi t}$ .
@@
The corresponding integral of $1/s$ equals $1$, so the above equality is still valid. Whenever $\lambda\in(0,\infty)$, $t\in[\varepsilon,1]$, $s$ is as in (8.12), and $\theta\lambda\leqslant|u_{k}|\leqslant\lambda$ for $1\leqslant k\leqslant m-1$, then we define
@@
$\qquad\qquad\qquad\qquad\times\Big(\int_{\mathbb{R}^{2}}\mathbbm{1}_{A}(x,y^{\prime})\mathbbm{h}^{(l)}_{s}(y^{\prime}-z^{\prime})\,\textup{d}y^{\prime}\Big)(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\,\textup{d}z\,\textup{d}z^{\prime}.$
@@
$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}.$
@@
$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}$
@@
$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}$ .
@@
$\times\int_{\varepsilon}^{1}\int_{\mathbb{R}}\int_{te^{-(m-1)\alpha-1}}^{te^{-(m-1)\alpha}}\int_{\mathbb{R}^{m-1}}\int_{\mathbb{R}^{2}}\big|\widehat{\mathbbm{1}_{A}}(x,\xi)\big|^{2}e^{-2\pi s^{2}|\xi|^{2}}|\xi|^{2}\,\textup{d}\xi\,\textup{d}x\,s\,\textup{d}s\,\textup{d}\alpha\,\frac{\textup{d}t}{t}.$
@@
where $C_{m}$ is the constant from Theorem 1. If we can find an axes-aligned cube $Q\subset\mathbb{R}^{n}$ of side-length $R_{0}$ on which $A$ has density greater than $(C_{m}/\log R_{0})^{1/(9m^{2})}$, then part (a) of that same theorem will apply to this cube $Q$, only translated from the position $0,R_{0}]^{n}$. It will give us a right $m$-simplex of volume 11 with vertices in $A\cap Q$. Moreover, the ratio of lengths of any two of its perpendicular edges will be at most
@@
Finally, for an arbitrary $V>0$ the set $V^{-1/m}A$ clearly has the same upper Banach density $\bar{\delta}_{n}(A)$, so the previous case applies to it and gives a right simplex of volume 11 with vertices in $V^{-1/m}A$ and with controlled ratios of lengths of its perpendicular edges. It remains to scale the whole picture back by factor $V^{1/m}$ and thus obtain the desired simplex of volume $V$. ∎